eviechocolate:

A video game that usually costs $30.65 is marked down 60%. Kelvin determined that the new price of the game would be $18.39. Look at Kelvin’s work and find his error. ($30.65)(0.60) = $18.39

jhonyy9:

use the rule of 3 simple and write 30,65 mean 100% x mean 60% x = 30,65 *60/100 = 18,39 so this is the sum with the prize was marked down so in this case the new prize will be 30,65 - 18,39 = ?$ hope helped understandably easy